OpenStudy (anonymous):

How many 4-digit positive integers can be made using the digits 0,1,2,3,4,5 and repetitions of digits are not allowed?

OpenStudy (anonymous):

4096

OpenStudy (anonymous):

6*6*6*6 = 6^4

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

you might have to remove some trivial cases

OpenStudy (anonymous):

the answer at the book is 300 but i don't know how to arrive at the answer

OpenStudy (anonymous):

ok it is 5*6*6*6

OpenStudy (anonymous):

oh repetition not allowed

OpenStudy (anonymous):

so 5*5*4*3

OpenStudy (anonymous):

so thats 300

OpenStudy (anonymous):

why is the second number 5

OpenStudy (anonymous):

ok let me explain the logic

OpenStudy (anonymous):

the first number in the thousandth's place is to choose from { 1,2,3,4,5} , for the hundreth place you can choose 0 , so you again have 5 choices .

OpenStudy (anonymous):

so for example , if you choose 1 for thousandth place, then you have { 0,2,3,4,5} numbers to choose for hundredth place.

OpenStudy (anonymous):

ooohhh that make sense. I thought it was 5,6,4,3 thanks a lot

OpenStudy (anonymous):

if you choose 3 for hundreth place, then you have { 0,2,4,5} for tens place

OpenStudy (anonymous):

5643?

OpenStudy (anonymous):

oh 5*6*4*3

OpenStudy (anonymous):

you're right, it is 5,5,4,3

OpenStudy (anonymous):

the trick is you have to add zero to the second choice, third, etc

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

the zero does not apply to the first choice. so its tricky ;)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

also to be clear they should say that 0430 is not a 4 digit number for example

OpenStudy (anonymous):

also interesting you have to do the choices in order, you cant skip around

OpenStudy (anonymous):

this is called multiplication principle by the way

OpenStudy (anonymous):

jm, any questions?

OpenStudy (anonymous):

5*5*4*3 = 300 Cannot use 0 for the thousands digit, therefore, one of the 5 non-zero digits can be chosen. For the hundreds digit, 0 and 4 of the other numbers can be chosen. The tens digit 4 and the units digit 3.